What is the angle between the hour and the minute hands if the time is 8:20?
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

What is the angle between the hour and the minute hands if the time is 8:20?
-------------------------
The minute hand is 1/3 of a revolution from straight up, or 120º
The hour hand makes one rev in 12 hours, so it have rotated 360º*((8 1/3)/12) degrees.
360º*((8 1/3)/12) = 30*(8 1/3) = 250º
250 - 120 = 130º

Answer by stanbon(75887)

What is the angle between the hour and the minute hands if the time is 8:20?
----------------
Draw the picture. There are 12 segments around the clock. Each segment
has 360/12 = 30 degrees.
If you assume the hour hand stays on 8 even when it is 8:20
the angle is 4*30 degrees = 120 degrees.
-----------
If you assume the hour hand moves (1/3) of the segment as the minute hand
moves to 20 after the hour, the angle is 120+(1/3)30 = 130 degrees.
------------------
Cheers,
Stan H.